OpenStudy (anonymous):

Help! I dont know how to do polar coordinates. Find all polar coordinates of point P where P = ordered pair (3, negative pi divided by 3).

OpenStudy (phi):

if you add 360 degrees (or 2pi radians) to the angle, you will get the same point in fact if you add *any* integer multiple of 2pi you get the same angle. people show this by writing \[ \left(3, -\frac{\pi}{3} + 2 \pi n \right) \text{ , n an integer}\]

OpenStudy (is3535):

can someone help

OpenStudy (phi):

if you allow negative distances, then you can write -3 for the distance, and compensate by add pi (180 degrees). And also any multiple of 2 pi, thus \[ \left(-3, -\frac{\pi}{3} + 3 \pi n \right) \text{ , n an integer} \]

OpenStudy (is3535):

@phi can you help me

OpenStudy (phi):

\[ \left(-3, -\frac{\pi}{3} + \pi+2 \pi n \right) \text{ , n an integer} \]

OpenStudy (anonymous):

\[(3,\frac{- \pi }{ 3 }+2n \pi )=\left( r, \theta \right)\],where n is an integer